# Multilevel Control Functional

### v3

@@ 1 Introduction
where $\Pi$ is a distribution with a Lebesgue density $\pi$ on $\mathcal{X}\subseteq\mathbb{R}^{d}$, and $f:\mathcal{X}\to\mathbb{R}$ is the integrand of interest. Assume that $f$ is square-integrable, that is, $\Pi[f^{2}]<\infty$. This is a common challenge in diverse applied fields such as finance (Glasserman, 2004; Chen et al., 2024), aerospace engineering (Morio and Balesdent, 2015; Geraci et al., 2017), hazard analysis (Geist and Parsons, 2006; Dalbey et al., 2008), medical physics (Rogers, 2006), among many others. This also frequently arises in statistics and machine learning, such as computing normalizing constants in probabilistic models (Atay-Kayis and Massam, 2005; Ohsaka and Matsuoka, 2020; Chehab et al., 2023), performing Bayesian inference (Golinski et al., 2019), training energy-based models (Song and Kingma, 2021) and variational inference (Buchholz et al., 2018; Vahdat and Kautz, 2020; Fujisawa and Sato, 2021).
@@
Monte Carlo (MC) (Liu, 2001; Rubinstein and Kroese, 2016) is the most widely used approach for estimating the integrals defined in Equation˜1. However, MC estimators tend to have high variance and slow convergence rates (Assaraf and Caffarel, 1999; Oates et al., 2019). Meanwhile, when dealing with complex scientific models, sampling or evaluating the integrand can be computationally expensive, e.g., large-scale computer simulators or costly experiments (Sánchez-Linares et al., 2016). To achieve the desired accuracy, the overall sampling and evaluation cost with standard MC can be prohibitive. One way to improve the efficiency of MC estimators is to reduce the variance of the integrand by control variates. Control variates (CVs) (Robert et al., 1999), including both parametrized CVs (Assaraf and Caffarel, 1999; South et al., 2022; Sun et al., 2023a) and non-parameterized CVs (Oates et al., 2017, 2019), are variance reduction techniques for MC estimators. This is achieved by designing and learning a function $g$ (known as control variate) well correlated to the integrand $f$. Another approach for computationally expensive functions is the use of multi-fidelity methods (Peherstorfer et al., 2018; Fernández-Godino, 2023), which have shown to be a useful scheme by compensating high-cost function approximations with low-cost ones, thereby reducing the overall computational cost. This particular scheme is highly related to the framework of multilevel Monte Carlo (MLMC) (Giles, 2008, 2015), which is designed for expensive integrands when cheaper approximations are available or can be constructed at several levels, e.g. tsunami modeling (Sánchez-Linares et al., 2016; Li et al., 2025). MLMC uses a sequence of multi-fidelity models to construct a telescoping sum for the original integral in Equation˜1. The telescoping sum consists of the expectations of increments between successive multi-fidelity models, namely $f_{l}$ and $f_{l-1}$, with $l$ indexing the hierarchy of accuracy levels. Given a fixed computational budget, MLMC yields more accurate estimates than standard MC estimators (Giles, 2015). This gain in performance has led to the widespread use of MLMC, not only in scientific applications but also as a powerful tool for enhancing many areas in machine learning, such as variational inference (Fujisawa and Sato, 2021). From the perspective of variance reduction for MC, MLMC treats the low-fidelity model $f_{l-1}$ as a control variate for the high-fidelity model $f_{l}$. This implies, however, that the variance of their differences is not minimized. Meanwhile, the existing work of CVs does not fully explore and exploit the inherent property of these multi-fidelity models of which the ‘precision’ can be decided and altered. See Figure˜1 for an illustrative example, with the corresponding results in Section˜D.1.
@@
Motivated by the above insights and unsolved problems, we propose a broadly applicable extension of control variates that integrates the strengths of control variates and multi-fidelity methods, in a manner analogous to multilevel Monte Carlo (Giles, 2008, 2015). In particular, we focus on non-parametric Stein-based control variates, control functionals (CFs), which naturally lead to a novel class of estimators named Multilevel Control Functionals (MLCFs). By leveraging the multi-fidelity structure alongside a derived variance bound allows us to assign the optimal sample sizes across all fidelity levels. Meanwhile, as we shown in Section˜3 and Section˜4, MLCFs are widely applicable. It can be applied to inference problems under un-normalized densities which are common in Bayesian inference, and can also be extended to variational inference as shown in this work.
@@
Although several related methods exist, their practical use is often constrained by inherent limitations. For example, (Li et al., 2023) combined Bayesian quadrature with MLMC, but the method was restricted to specific kernel-distribution pairs and was not suitable for complex distributions such as unnormalized distributions in Bayesian inference. Similarly, existing works on multilevel control variates are tailored for specific cases. Nobile and Tesei (2015) used the solution to an auxiliary diffusion problem with smoothed coefficients as the control variate, which was only applicable to specific partial differential equations. Fairbanks et al. (2017) used a low-rank representation for high-fidelity models to construct a control variate. Geraci et al. (2017) used a simplified physical model, which required additional expert knowledge. In contrast, the proposed MLCFs are broadly applicable and can be implemented without reliance on domain-specific expertise.
@@ 2.1 Control Variates
The above estimator follows a central limiting theorem: $\sqrt{n}(\hat{\Pi}_{\textup{MC}}[f]-\Pi[f])\xrightarrow[]{}\mathcal{N}(0,\mathbb{V}[f])$. Thus, the convergence rate of the estimator is often determined by the sample size $n$ and the variance $\mathbb{V}[f]$. It often requires a large number of function evaluations to achieve the desired accuracy. Similar results hold for Markov Chain Monte Carlo (MCMC) (Dellaportas and Kontoyiannis, 2012; Alexopoulos et al., 2023) and quasi-Monte Carlo methods (Hickernell et al., 2005). In this work, we will focus on the Monte Carlo case. One way to improve the performance of Monte Carlo (MC) estimators is to identify a function $g\in\mathcal{L}^{2}(\Pi)$ with known mean $\Pi[g]$ such that $\mathbb{V}[f-g]$ is small. Such $g$ is also known as control variates (CVs). Finding $g$ with a known mean $\Pi[g]$ is then the first challenge. When $\Pi$ is relatively simple, ad-hoc methods such as Taylor expansions of the integrand $f$ can be used (Paisley et al., 2012; Wang et al., 2013). While for more general and complex distributions, the ones often encountered in Bayesian inference, we can employ Stein’s method (Anastasiou et al., 2023) to construct such functions which are also known as Stein-based control variates.
@@
Stein-based control variates (Oates et al., 2017; Si et al., 2022; South et al., 2022; Sun et al., 2023b) are variance reduction tools for Monte Carlo integration. They are also widely used in the cases when the density is unnormalized and when the samples are MCMC samples, which often appears in Bayesian inference. The first step is to construct a candidate set $\mathcal{G}$ such that $\Pi[g]=0$ for $\forall g\in\mathcal{G}$. This can be achieved by using Stein’s operators $\mathcal{S}_{\Pi}$ (e.g. the Langevin Stein operator); see (Anastasiou et al., 2023) for a detailed review. By using the zero-mean property, we have $\Pi[f-g]=\Pi[f]$ for $\forall g\in\mathcal{G}$. The second step is to select an effective control variate $g\in\mathcal{G}$ with reduced variance, i.e., $\mathbb{V}[f-g]=\Pi[(f-g-\Pi[f-g])^{2}]<\mathbb{V}[f]$ (Oates et al., 2017; Zhu et al., 2019; South et al., 2022). Such an effective control variate $g$ is often learnt by minimizing the empirical (penalized) variance of $\mathbb{V}[f-g]$ conditioning on $m$ samples $\{x_{i}\}_{i=1}^{m}$ and their function evaluations from all samples $\{x_{i}\}_{i=1}^{n}$ available. Then, through estimating $\Pi[f-g]$ with the remaining nmn-m function evaluations, we can get an estimate of $\Pi[f]$ with reduced variance and improved accuracy, given by
@@
We consider a non-parametric family of control variates, control functionals (CFs) (Oates et al., 2017, 2019), which is designed for single Monte Carlo integration problems. It is a class of non-parametric Stein-based control variates based on reproducing kernel Hilbert spaces (RKHS). It applies the Langevin Stein operator $\mathcal{S}_{\Pi}$ onto vector-valued functions $u\in C^{1}(\mathcal{X})\times\cdots\times C^{1}(\mathcal{X})$ which takes the form $\mathcal{S}_{\Pi}[u](x):=\nabla_{x}\cdot u(x)+u(x)\cdot\nabla_{x}\log\pi(x)$ where $\nabla\cdot$ denotes the divergence operator and $\nabla$ denotes the gradient operator. Each component function $u_{i}:\mathcal{X}\to\mathbb{R}$ is constrained to belong to a Hilbert space $\mathcal{H}$. Let $\mathcal{H}_{k}$ be the RKHS induced by a reproducing kernel $k$. The image of $\mathcal{U}:=\mathcal{H}_{k}\times\cdots\times\mathcal{H}_{k}$ under $\mathcal{S}_{\Pi}$ is a RKHS $\mathcal{G}$ with kernel $k_{0}$ (also known as a Stein kernel); see Section˜A.2 in Appendix A.2. Oates et al. (2017, 2019) used functional approximations $s(x)=\beta+\mathcal{S}_{\Pi}[u](x)$ where $\beta$ and $u$ are selected by solving a constraint least-square optimisation problem in $\mathcal{G}$ conditioning on $m$ samples $\{x_{i}\}_{i=1}^{m}$ and $\{f(x_{i})\}_{i=1}^{m}$. The control functional takes the form of: $g_{m}(x)=s(x)-\Pi[s]$. The standard control functional estimator is given by
@@
where $X^{0}=(x_{1},\ldots,x_{m})^{\top}$, $X^{1}=(x_{m+1},\ldots,x_{n})^{\top}$, $(f(X^{0}))_{i}=f(x_{i})$, $(k_{0}(X^{0},X^{0}))_{i,j}=k_{0}(x_{i},x_{j})$, for all $i,j\in\{1,\ldots,m\}$, and $(f(X^{1}))_{i}=f(x_{m+i})$, $(k_{0}(X^{1},X^{0}))_{i,j}=k_{0}(x_{m+i},x_{j})$, for all $i\in\{1,\ldots,n-m\}$, and for all $j\in\{1,\ldots,m\}$. A major drawback of control functional is the $\mathcal{O}(m^{3})$ computational cost, which can be alleviated by using stochastic optimization as in (Zhu et al., 2019; Si et al., 2022; Sun et al., 2023b, a). Meanwhile, this would not be an severe issue in the setting considered in this work as such cost is much smaller than the cost of the evaluation of integrand.
@@ 2.2 Multi-fidelity Models and Multilevel Monte Carlo
Multilevel Monte Carlo (Giles, 2008, 2015) uses a hierarchy of approximations $f_{0},f_{1},\ldots,f_{L-1}$ to $f_{L}:=f$ with increasing levels of accuracy and cost to estimate the integral of interest. The method can achieve a higher accuracy with a lower computational cost compared to MC using only the $f_{L}:=f$. Given the sequence of approximations, MLMC sums up the estimates of the corrections with respect to the consecutive lower level and obtain the telescoping sum
@@
From the view of variance reduction, $f_{l-1}$ can be regarded as a control variate for $f_{l}$ for all levels. It has shown that by carefully analysis of the number of samples assigned to each level, MLMC can largely improve efficiency over standard MC when the desired accuracy is fixed (Giles, 2015). Such MLMC methods have also shown successes in various fields including probabilistic numeric (Li et al., 2023), simulation-based inference for multiple simulators of various fidelity (Hikida et al., 2025) and variational inference (Fujisawa and Sato, 2021; Shi and Cornish, 2021).
@@ 3 Methodology
The result is then used to compute the optimal sample size at each level for the synthetic example in Section˜4. Since the constant $r_{l}$ is independent of $f_{l}$, $f_{l-1}$ and data points, and the domain, $d$ and $\tau$ are the same across all levels, the effect of $r_{l}$ could be normalized away. Alternatively, one may assume a uniform bound $r\geq\max_{l=0,1,\dots,L}r_{l}$ and use this in the proof of the theorem to get a result without $r_{l}$. The RKHS norm can be computed using reproducing property of reproducing kernels for the synthetic example. Formally, Kanagawa et al. (2018, Theorem 2.4) provides a general equation for the RKHS norm. In practical applications, though functions may not exist in the form of linear combinations of kernel functions, users can use data-based methods to estimate the RKHS norm, For example, Scharnhorst et al. (2022, Appendix A) provides an example of estimating the RKHS norm using randomly sampled data. There are also other data-driven approaches illustrated in Karvonen (2022); Tokmak et al. (2025a, b).
@@
MLCFs can be further extended for variational inference (VI). To be precise, we consider the scenarios when the objective is to minimize the Kullback–Leibler (KL) divergence between the variational distribution $q(z|\lambda)$ and the posterior $p(z|D)$ ($D$ denotes the observations) with respect to the parameters $\lambda$. This minimization is equivalent to maximizing the evidence lower bound (ELBO): $\mathcal{L}(\lambda)=\mathbb{E}_{q(z|\lambda)}[\log p(z,D)-\log q(z|\lambda)]$. We will focus on the re-parameterized gradient estimator to optimize the objective function, where $z=\mathcal{T}(x,\lambda)$ is expressed as a deterministic transformation $\mathcal{T}$ of a noise variable $x$ with distribution $\pi(x)$. Fujisawa and Sato (2021) proposed the multilevel re-parameterized gradient (MLRG), which reduces the variance by recycling the previous parameters and gradients. At iteration $L$, the multilevel re-parameterized gradient (MLRG) has the following form: $\nabla_{\lambda_{L}}^{\text{MLRG}}\mathcal{L}(\lambda_{L})=\sum_{l=0}^{L}\Pi[f_{\lambda_{l}}-f_{\lambda_{l-1}}]$ where $f_{\lambda_{l}}(x)=\nabla_{\lambda_{l}}\log p(D,\mathcal{T}(x,\lambda_{l}))-\nabla_{\lambda_{l}}\log q(\mathcal{T}(x,\lambda_{l})|\lambda_{l})$ with $f_{\lambda_{-1}}(x)=0$ for notational convenience. Fujisawa and Sato (2021) used Monte Carlo to estimate MLRG and resulted in multilevel Monte Carlo re-parameterized gradient (MLMCRG) estimators, which took the form of $\hat{\nabla}_{\lambda_{L}}^{\text{MLRG}}\mathcal{L}(\lambda_{L})=\hat{\Pi}_{\textup{MLMC}}[f_{\lambda_{L}}]:=\sum_{l=0}^{L}\hat{\Pi}_{\textup{MC}}[f_{\lambda_{l}}-f_{\lambda_{l-1}}]$. The proposed MLCF then naturally results in a novel series of estimators, named as multilevel control functional re-parameterized gradient (MLCFRG) estimators:

### v4

@@ 1 Introduction
where $\Pi$ is a distribution with a Lebesgue density $\pi$ on $\mathcal{X}\subseteq\mathbb{R}^{d}$, and $f:\mathcal{X}\to\mathbb{R}$ is the integrand of interest. Assume that $f$ is square-integrable, that is, $\Pi[f^{2}]<\infty$. This is a common challenge in diverse applied fields such as finance (Glasserman, 2004; Chen et al., 2024), aerospace engineering (Morio & Balesdent, 2015; Geraci et al., 2017), hazard analysis (Geist & Parsons, 2006; Dalbey et al., 2008), medical physics (Rogers, 2006), among many others. This also frequently arises in statistics and machine learning, such as computing normalizing constants in probabilistic models (Atay-Kayis & Massam, 2005; Ohsaka & Matsuoka, 2020; Chehab et al., 2023), performing Bayesian inference (Golinski et al., 2019), training energy-based models (Song & Kingma, 2021) and variational inference (Buchholz et al., 2018; Vahdat & Kautz, 2020; Fujisawa & Sato, 2021).
@@
Monte Carlo (MC) (Liu, 2001; Rubinstein & Kroese, 2016) is the most widely used approach for estimating the integrals defined in Equation˜1. However, MC estimators tend to have high variance and slow convergence rates (Assaraf & Caffarel, 1999; Oates et al., 2019). Meanwhile, when dealing with complex scientific models, sampling or evaluating the integrand can be computationally expensive, e.g., large-scale computer simulators or costly experiments (Sánchez-Linares et al., 2016). To achieve the desired accuracy, the overall sampling and evaluation cost with standard MC can be prohibitive. One way to improve the efficiency of MC estimators is to reduce the variance of the integrand by control variates. Control variates (CVs) (Robert et al., 1999), including both parametrized CVs (Assaraf & Caffarel, 1999; South et al., 2022; Sun et al., 2023b; Sun, 2023) and non-parameterized CVs (Oates et al., 2017; 2019), are variance reduction techniques for MC estimators. This is achieved by designing and learning a function $g$ (known as control variate) well correlated to the integrand $f$. Another approach for computationally expensive functions is the use of multi-fidelity methods (Peherstorfer et al., 2018; Fernández-Godino, 2023), which have shown to be a useful scheme by compensating high-cost function approximations with low-cost ones, thereby reducing the overall computational cost. This particular scheme is highly related to the framework of multilevel Monte Carlo (MLMC) (Giles, 2008; 2015), which is designed for expensive integrands when cheaper approximations are available or can be constructed at several levels, e.g. tsunami modeling (Sánchez-Linares et al., 2016; Li, 2024; Li et al., 2025). MLMC uses a sequence of multi-fidelity models to construct a telescoping sum for the original integral in Equation˜1. The telescoping sum consists of the expectations of increments between successive multi-fidelity models, namely $f_{l}$ and $f_{l-1}$, with $l$ indexing the hierarchy of accuracy levels. Given a fixed computational budget, MLMC yields more accurate estimates than standard MC estimators (Giles, 2015). This gain in performance has led to the widespread use of MLMC, not only in scientific applications but also as a powerful tool for enhancing many areas in machine learning, such as variational inference (Fujisawa & Sato, 2021). From the perspective of variance reduction for MC, MLMC treats the low-fidelity model $f_{l-1}$ as a control variate for the high-fidelity model $f_{l}$. This implies, however, that the variance of their differences is not minimized. Meanwhile, the existing work of CVs does not fully explore and exploit the inherent property of these multi-fidelity models of which the ‘precision’ can be decided and altered. See Figure˜1 for an illustrative example, with the corresponding results in Section˜D.1.
@@
Motivated by the above insights and unsolved problems, we propose a broadly applicable extension of control variates that integrates the strengths of control variates and multi-fidelity methods, in a manner analogous to multilevel Monte Carlo (Giles, 2008; 2015). In particular, we focus on non-parametric Stein-based control variates, control functionals (CFs), which naturally lead to a novel class of estimators named Multilevel Control Functionals (MLCFs). By leveraging the multi-fidelity structure alongside a derived variance bound allows us to assign the optimal sample sizes across all fidelity levels. Meanwhile, as we shown in Section˜3 and Section˜4, MLCFs are widely applicable. It can be applied to inference problems under un-normalized densities which are common in Bayesian inference, and can also be extended to variational inference as shown in this work.
@@
Although several related methods exist, their practical use is often constrained by inherent limitations. For example, (Li et al., 2023) combined Bayesian quadrature with MLMC, but the method was restricted to specific kernel-distribution pairs and was not suitable for complex distributions such as unnormalized distributions in Bayesian inference. Similarly, existing works on multilevel control variates are tailored for specific cases. Nobile & Tesei (2015) used the solution to an auxiliary diffusion problem with smoothed coefficients as the control variate, which was only applicable to specific partial differential equations. Fairbanks et al. (2017) used a low-rank representation for high-fidelity models to construct a control variate. Geraci et al. (2017) used a simplified physical model, which required additional expert knowledge. In contrast, the proposed MLCFs are broadly applicable and can be implemented without reliance on domain-specific expertise.
@@ 2.1 Control Variates
The above estimator follows a central limiting theorem: $\sqrt{n}(\hat{\Pi}_{\textup{MC}}[f]-\Pi[f])\xrightarrow[]{}\mathcal{N}(0,\mathbb{V}[f])$. Thus, the convergence rate of the estimator is often determined by the sample size $n$ and the variance $\mathbb{V}[f]$. It often requires a large number of function evaluations to achieve the desired accuracy. Similar results hold for Markov Chain Monte Carlo (MCMC) (Dellaportas & Kontoyiannis, 2012; Alexopoulos et al., 2023) and quasi-Monte Carlo methods (Hickernell et al., 2005). In this work, we will focus on the Monte Carlo case. One way to improve the performance of Monte Carlo (MC) estimators is to identify a function $g\in\mathcal{L}^{2}(\Pi)$ with known mean $\Pi[g]$ such that $\mathbb{V}[f-g]$ is small. Such $g$ is also known as control variates (CVs). Finding $g$ with a known mean $\Pi[g]$ is then the first challenge. When $\Pi$ is relatively simple, ad-hoc methods such as Taylor expansions of the integrand $f$ can be used (Paisley et al., 2012; Wang et al., 2013). While for more general and complex distributions, the ones often encountered in Bayesian inference, we can employ Stein’s method (Anastasiou et al., 2023) to construct such functions which are also known as Stein-based control variates.
@@
Stein-based control variates (Oates et al., 2017; Si et al., 2022; South et al., 2022; Sun et al., 2023a; b) are variance reduction tools for Monte Carlo integration. They are also widely used in the cases when the density is unnormalized and when the samples are MCMC samples, which often appears in Bayesian inference. The first step is to construct a candidate set $\mathcal{G}$ such that $\Pi[g]=0$ for $\forall g\in\mathcal{G}$. This can be achieved by using Stein’s operators $\mathcal{S}_{\Pi}$ (e.g. the Langevin Stein operator); see (Anastasiou et al., 2023) for a detailed review. By using the zero-mean property, we have $\Pi[f-g]=\Pi[f]$ for $\forall g\in\mathcal{G}$. The second step is to select an effective control variate $g\in\mathcal{G}$ with reduced variance, i.e., $\mathbb{V}[f-g]=\Pi[(f-g-\Pi[f-g])^{2}]<\mathbb{V}[f]$ (Oates et al., 2017; Zhu et al., 2019; South et al., 2022). Such an effective control variate $g$ is often learnt by minimizing the empirical (penalized) variance of $\mathbb{V}[f-g]$ conditioning on $m$ samples $\{x_{i}\}_{i=1}^{m}$ and their function evaluations from all samples $\{x_{i}\}_{i=1}^{n}$ available. Then, through estimating $\Pi[f-g]$ with the remaining nmn-m function evaluations, we can get an estimate of $\Pi[f]$ with reduced variance and improved accuracy, given by
@@
We consider a non-parametric family of control variates, control functionals (CFs) (Oates et al., 2017; 2019), which is designed for single Monte Carlo integration problems. It is a class of non-parametric Stein-based control variates based on reproducing kernel Hilbert spaces (RKHS). It applies the Langevin Stein operator $\mathcal{S}_{\Pi}$ onto vector-valued functions $u\in C^{1}(\mathcal{X})\times\cdots\times C^{1}(\mathcal{X})$ which takes the form $\mathcal{S}_{\Pi}[u](x):=\nabla_{x}\cdot u(x)+u(x)\cdot\nabla_{x}\log\pi(x)$ where $\nabla\cdot$ denotes the divergence operator and $\nabla$ denotes the gradient operator. Each component function $u_{i}:\mathcal{X}\to\mathbb{R}$ is constrained to belong to a Hilbert space $\mathcal{H}$. Let $\mathcal{H}_{k}$ be the RKHS induced by a reproducing kernel $k$. The image of $\mathcal{U}:=\mathcal{H}_{k}\times\cdots\times\mathcal{H}_{k}$ under $\mathcal{S}_{\Pi}$ is a RKHS $\mathcal{G}$ with kernel $k_{0}$ (also known as a Stein kernel); see Section˜A.2 in Appendix A.2. Oates et al. (2017; 2019) used functional approximations $s(x)=\beta+\mathcal{S}_{\Pi}[u](x)$ where $\beta$ and $u$ are selected by solving a constraint least-square optimisation problem in $\mathcal{G}$ conditioning on $m$ samples $\{x_{i}\}_{i=1}^{m}$ and $\{f(x_{i})\}_{i=1}^{m}$. The control functional takes the form of: $g_{m}(x)=s(x)-\Pi[s]$. The standard control functional estimator is given by
@@
where $X^{0}=(x_{1},\ldots,x_{m})^{\top}$, $X^{1}=(x_{m+1},\ldots,x_{n})^{\top}$, $(f(X^{0}))_{i}=f(x_{i})$, $(k_{0}(X^{0},X^{0}))_{i,j}=k_{0}(x_{i},x_{j})$, for all $i,j\in\{1,\ldots,m\}$, and $(f(X^{1}))_{i}=f(x_{m+i})$, $(k_{0}(X^{1},X^{0}))_{i,j}=k_{0}(x_{m+i},x_{j})$, for all $i\in\{1,\ldots,n-m\}$, and for all $j\in\{1,\ldots,m\}$. A major drawback of control functional is the $\mathcal{O}(m^{3})$ computational cost, which can be alleviated by using stochastic optimization as in (Zhu et al., 2019; Si et al., 2022; Sun et al., 2023a; b). Meanwhile, this would not be an severe issue in the setting considered in this work as such cost is much smaller than the cost of the evaluation of integrand.
@@ 2.2 Multi-fidelity Models and Multilevel Monte Carlo
Multilevel Monte Carlo (Giles, 2008; 2015) uses a hierarchy of approximations $f_{0},f_{1},\ldots,f_{L-1}$ to $f_{L}:=f$ with increasing levels of accuracy and cost to estimate the integral of interest. The method can achieve a higher accuracy with a lower computational cost compared to MC using only the $f_{L}:=f$. Given the sequence of approximations, MLMC sums up the estimates of the corrections with respect to the consecutive lower level and obtain the telescoping sum
@@
From the view of variance reduction, $f_{l-1}$ can be regarded as a control variate for $f_{l}$ for all levels. It has shown that by carefully analysis of the number of samples assigned to each level, MLMC can largely improve efficiency over standard MC when the desired accuracy is fixed (Giles, 2015). Such MLMC methods have also shown successes in various fields including probabilistic numeric (Li et al., 2023), simulation-based inference for multiple simulators of various fidelity (Hikida et al., 2025) and variational inference (Fujisawa & Sato, 2021; Shi & Cornish, 2021).
@@ 3 Methodology
The result is then used to compute the optimal sample size at each level for the synthetic example in Section˜4. Since the constant $r_{l}$ is independent of $f_{l}$, $f_{l-1}$ and data points, and the domain, $d$ and $\tau$ are the same across all levels, the effect of $r_{l}$ could be normalized away. Alternatively, one may assume a uniform bound $r\geq\max_{l=0,1,\dots,L}r_{l}$ and use this in the proof of the theorem to get a result without $r_{l}$. The RKHS norm can be computed using reproducing property of reproducing kernels for the synthetic example. Formally, Kanagawa et al. (2018, Theorem 2.4) provides a general equation for the RKHS norm. In practical applications, though functions may not exist in the form of linear combinations of kernel functions, users can use data-based methods to estimate the RKHS norm, For example, Scharnhorst et al. (2022, Appendix A) provides an example of estimating the RKHS norm using randomly sampled data. There are also other data-driven approaches illustrated in Karvonen (2022); Tokmak et al. (2025a; b).
@@
MLCFs can be further extended for variational inference (VI). To be precise, we consider the scenarios when the objective is to minimize the Kullback–Leibler (KL) divergence between the variational distribution $q(z|\lambda)$ and the posterior $p(z|D)$ ($D$ denotes the observations) with respect to the parameters $\lambda$. This minimization is equivalent to maximizing the evidence lower bound (ELBO): $\mathcal{L}(\lambda)=\mathbb{E}_{q(z|\lambda)}[\log p(z,D)-\log q(z|\lambda)]$. We will focus on the re-parameterized gradient estimator to optimize the objective function, where $z=\mathcal{T}(x,\lambda)$ is expressed as a deterministic transformation $\mathcal{T}$ of a noise variable $x$ with distribution $\pi(x)$. Fujisawa & Sato (2021) proposed the multilevel re-parameterized gradient (MLRG), which reduces the variance by recycling the previous parameters and gradients. At iteration $L$, the multilevel re-parameterized gradient (MLRG) has the following form: $\nabla_{\lambda_{L}}^{\text{MLRG}}\mathcal{L}(\lambda_{L})=\sum_{l=0}^{L}\Pi[f_{\lambda_{l}}-f_{\lambda_{l-1}}]$ where $f_{\lambda_{l}}(x)=\nabla_{\lambda_{l}}\log p(D,\mathcal{T}(x,\lambda_{l}))-\nabla_{\lambda_{l}}\log q(\mathcal{T}(x,\lambda_{l})|\lambda_{l})$ with $f_{\lambda_{-1}}(x)=0$ for notational convenience. Fujisawa & Sato (2021) used Monte Carlo to estimate MLRG and resulted in multilevel Monte Carlo re-parameterized gradient (MLMCRG) estimators, which took the form of $\hat{\nabla}_{\lambda_{L}}^{\text{MLRG}}\mathcal{L}(\lambda_{L})=\hat{\Pi}_{\textup{MLMC}}[f_{\lambda_{L}}]:=\sum_{l=0}^{L}\hat{\Pi}_{\textup{MC}}[f_{\lambda_{l}}-f_{\lambda_{l-1}}]$. The proposed MLCF then naturally results in a novel series of estimators, named as multilevel control functional re-parameterized gradient (MLCFRG) estimators: